Evaluate the function at the indicated values. (If an answer is undefined, enter UNDEFINED.)
f(x) = x3 + 8x; f(−2), f(−1), f(0), f(1/2)
If I am correct, f(0)=0


Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52794)

.

f(0) = 0.


It is correct.

You are correct.



Answer by math_tutor2020(3817)


Replace x with each input mentioned.
Use the order of operations (PEMDAS) to simplify.
f(x) = x^3+8x
f(-2) = (-2)^3+8(-2)
f(-2) = -8+8(-2)
f(-2) = -8-16
f(-2) = -24

Similar steps are followed for x = -1, x = 0, and x = 1/2
You are correct in stating that f(0) = 0.

The Desmos calculator can be used to verify.
Type f(x) = x^3+8x into one input box.
Then type f(-2) through f(1/2) to evaluate each input.
See this link for more info:
https://www.desmos.com/calculator/rorsajdy6w
Focus on the calculations in the left-side panel.